solve each system by graphing both lines and identify where they intersect:
y = -3x + 4
6y + 18x = 24
question 2
y = x + 3
x - 7 = 0

Answer by mananth(16946)

y = -3x + 4.................1
6y + 18x = 24
6y=-18x+24
/6
y=-3x+4......................2
Both the lines are having same slope -3 and same y intercept +4
so the lines coincide. ( they are one and the same lines)

when x= 0 y= 4 (0,4)
when y = 0 x= 1.33(1.33,0)

All the points on the line are the solutions to both the equations. So, the system of equations has infinite number of solutions.
So, the system of equations is consistent and the equations are dependent.